Prime takes a trip in spaceship to a distant station (distance L = 1 light year) at uniform speed (v=0.8c). 

After reaching the distant station, Prime takes a return spaceship to meet Unprime  in Ahmedabad. Explain who is younger after this journey: Prime or Unprime, and  by how many years? Use units of distance as light year and that of time as year.


1
Expert's answer
2021-01-06T13:57:18-0500

Explanations & Calculations

  • Time dilation/expansion is the base theory applied here.
  • And what this question describes is the thought called "twin paradox".
  • The earthbound Unprime sees that Prime takes a longer time to return to earth equally Prime sees that Unprime has aged more as he returns.
  • The relationship for the time dilation is given by,

"\qquad\qquad
\begin{aligned}
\small \Delta T&= \small \gamma\Delta \tau
\end{aligned}" "\qquad\qquad
\begin{aligned}
\small \tau&= \small \text{Lorentz factor}=\sqrt{1-\frac{v^2}{c^2}}\\
\small \Delta T&= \small \text{duration relative to earthbound one} \\
\small \Delta \tau &= \small \text{duration relative spacebound one}
\end{aligned}"

  • As seen by Prime, it takes "\small 2\times \frac{1c\times year}{0.8c}=2.5 years" for the whole trip.
  • Therefore, this time is seen as

"\qquad\qquad
\begin{aligned}
\small \Delta t&= \small \frac{1}{\sqrt{1-\frac{(0.8c)^2}{c^2}}}\times 2.5\text{years}\\
&=\small 4.167 \text{years}
\end{aligned}"

by Unprime.

  • Therefore, Prime is 4.167 years younger than Unprime.